OpenStudy (anonymous):

Challenge: Graph the equation with a diameter that has endpoints at (-2, -2) and (4, 2). Label the center and at least four points on the circle. Write the equation of the circle.

OpenStudy (math&ing001):

First you draw the center. Should be \[O(\frac{ 4+(-2) }{ 2 },\frac{ 2+(-2) }{ 2 })= O(1, 0)\] Then, to find points that belong to the circle, your points should have the same distance to O as (-2,-2) and (4,2).